a rectangle's length is 8 cm more than three times its width. If the perimeter is 128 cm, find the length and the width
Answer by macston(5194)

.
W=width; L=length=3W+8cm
.
P=2(L+W)
128cm=2((3W+8cm)+W)
64cm=4W+8cm
56cm=4W
14cm=W ANSWER 1: The width is 14cm.
L=3W+8cm=3(14cm)+8cm=50cm ANSWER 2: The length is 50 cm.
.
CHECK:
.
P=2(L+W)
128cm=2(50cm+14cm)
128cm=2(64cm)
128cm=128cm